OpenStudy (anonymous):

What is the slope and y-intercept of the graph of 3x + 6y = 36? A. slope = -2; y-intercept = (0, 36) B. slope = -2; y-intercept = (0, 6) C. slope = 6; y-intercept = (0,/1/2 ) D. slope =/1/2 ; y-intercept = (0, 6)

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

slope is the "M"

OpenStudy (anonymous):

First you need to make your equation y=mx+b. m is the slope and b is the y-intercept.

OpenStudy (anonymous):

I think D is to read slope =-1/2 ; y-intercept = (0, 6) Then it is the answer